OpenStudy (anonymous):

0.5x+2 ≤1.3x-7 solution?

OpenStudy (anonymous):

Multiply each side by 10 5x+20<=13x-7 20<=13x-5x-7 20<=8x-7 27<=8x 27/8<=x x>=27/8

OpenStudy (anonymous):

Ah yes my mistake...good eye

OpenStudy (anonymous):

so is the answer still x >= 27/8??

OpenStudy (anonymous):

Would be x>=90/8, x>=45/4